OpenStudy (anonymous):

The length of a rectangle is 48 in and is dilated to 12 in. What is the scale factor?

OpenStudy (nurali):

dilation factor is always equal to the new side measure / old side measure 48/12

OpenStudy (anonymous):

12/48=1/4 so the scale factor is 1/4, another way to look at this is 48/12=4, however, because the original got smaller in size, the answer must be a fraction, otherwise, when you take 48*4 you will not get 12. On the other hand, when you take 48*1/4 you do get 12.